The width of a rectangular playground is 2x - 5 feet, and tthe length is 3x +9 feet. Write polynomial P(x) that represents the perimeter and then evaluate this perimeter polynomial if x is 4 feet.
This is my first class in Algebra, kindly help me to solve this word problem.

Answer by tutorcecilia(2152)

Use the formula for perimeter:
P=2[length + width}
P=2[(2x-5)+(3x+9)] [plug-in the values and simplify]
.
P=2(5x+4)
P=10x+8
.
P(4)=10(4)+8
P(4)=40+8
P(4)=48